# Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up

### v1

@@ 3. Formulate Logical Pseudocode:
3.3 Cognitive Preference Manager
Algorithm 2 Cognitive Preference Manager (CPM)
1:  $P_{task}\leftarrow P,P^{*}_{task}\leftarrow P^{*}$
3:  if $s\geq\delta$ then
5:     Aggregate relevant task-specific knowledge.
6:     return  $P_{final}$.
7:  else
8:     Unknown detected: Adapt and expand $P_{task}$.
10:     return  $P_{final}$.
11:  end if
Cognitive Preference Manager.
By efficiently evaluating cognitive results, CPM integrates alternative aggregation strategies for $\mathcal{M_{LLM}}$ based on $S_{cog}$ 111we use $S_{cog}$ to conduct reverse evaluation of GPT-4 in our experiments., as detailed in appendix A: (1) Solution logic aggregation for known tasks: $\mathcal{M_{LLM}}$ merges beneficial aspects from the original prompt $P$ with the LLM-taste prompt $P^{*}$ to create the final prompt $P_{\textit{final}}$. (2) Stylistic template aggregation for unknown $\mathcal{M_{LLM}}$ extracts a cognitive preference template $T$ from the incorrect context in the LLM-taste prompt, and integrates meta-cognitive elements from the original prompt $P$ into $T$ to construct the final prompt $P_{final}$.
$answer=\mathcal{M_{LLM}}(P_{final},problem)$ (8)
4 Experiments
4.1 Datasets and Tasks
To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al. (2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al. (2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al. (2024) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24. (2) BIG-Bench Suzgun et al. (2023); Srivastava et al. (2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting from BIG-Bench Hard Suzgun et al. (2023) and Checkmate-in-One from BIG-Bench suite Srivastava et al. (2023); (3) Python Puzzles Schuster et al. (2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels.;(4) Multilingual Grade School Math (MGSM) Shi et al. (2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al. (2021), featuring translations of a subset of examples into ten diverse languages.(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three specified words verbatim.
4.2 Baselines
4.3 Experiment Setup
To ensure fair comparisons with previous methods, we utilize GPT-4, the latest model supporting $logprobs$  [^2] https://platform.openai.com/docs/api-reference/introduction, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study. For the $warm$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API. Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency. For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^3] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al. (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
4.4 Evaluation Metrics
Task CoT (NeurIPS) Wei et al. (2022) Meta-prompting Suzgun and Kalai (2024) ToT (NeurIPS) Yao et al. (2024) GoT (AAAI) Besta et al. (2024) BoT (NeurIPS) Yang et al. (2024a) RoT (Ours)
1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
Game of 24 31.5 33.2 65.4 66.8 73.3 74.1 73.0 73.5 82.8 82.5 97.5 97.8
Geometric Shapes 57.3 60.8 76.8 78.3 58.8 60.1 56.1 57.8 90.1 90.9 94.8 95.0
Multi-Step Arithmetic 83.4 85.3 89.5 90.6 88.5 90.3 87.5 88.8 94.6 96.8 98.6 99.0
WordSorting 80.8 82.5 97.8 98.6 95.0 95.8 95.6 96.8 99.2 99.6 99.8 100.0
Checkmate-in-One 35.8 36.5 58.0 57.5 49.8 48.8 49.0 51.2 87.2 88.1 91.8 92.0
Python Puzzles 35.3 35.8 45.0 45.5 43.8 45.1 42.5 43.5 51.8 52.6 54.8 57.5
MGSM (avg) 83.8 84.2 85.2 85.4 84.5 86.1 85.8 86.8 87.5 87.3 88.9 90.1
Sonnet Writing 66.5 70.2 78.8 79.8 68.9 69.8 68.0 62.5 79.0 79.5 89.5 92.0
Knowledge Boundary and Cognitive Preference Consistency.
As described in Section 3.2, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$. We first evaluate GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^4] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot. and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
Reasoning Accuracy.
Following Suzgun and Kalai (2024); Yao et al. (2024), we introduce a LLM555we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong). We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
Reasoning Efficiency.
We evaluate reasoning efficiency by calculating ${t_{eff}}$, calculated as the average time spent per task across all samples in the test dataset:
$t_{eff}=\frac{1}{N}\sum_{j=1}^{M}\sum_{i=1}^{N}t_{ij}$ (9)
5 Results and Discussion
Task 1-shot 2-shot
Kno. Unkno. Kno. Unkno.
Game of 24
MGSM (avg)
Multi-Step Arithmetic
WordSorting
Python Puzzles
Geometric Shapes
Checkmate-in-One
Sonnet Writing
Table 2: Knowledge Boundary under Reversal Reasoning. Kno. for Known, Unkno. for Unknown.
5.1 Knowledge Boundary and Cognitive Preference Consistency
Knowledge Boundary.
As shown in Table 2, our findings provide preliminary support for our hypothesis concerning cognitive knowledge boundaries of LLMs in reverse reasoning across current task benchmarks in one-shot and two-shot settings Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024). Notably, GPT-4 excels in structured reasoning tasks such as Game of 24, Geometric Shapes, and Checkmate-in-One, consistently achieving strong results in both one-shot and two-shot settings. In contrast, tasks such as MGSM (avg) and Python Puzzles fall into an unknown category in the one-shot scenario, stemming from multi-source problems that are challenging to make reverse reasoning without sufficient context, indirectly supporting the rationale for our CPM approach, detailed in appendix B.
Cognitive Preference Consistency.
Moreover, our human evaluation of $\textit{Con}_{cog}$ produced a high score of 4.32 in the two-shot setting and 4.01 in the one-shot setting, further validating that LLMs exhibit cognitive preferences shaped by RLHF Ouyang et al. (2022).
5.2 Reasoning Accuracy and Efficiency
Reasoning Accuracy.
As shown table 1, our RoT model consistently outperforms all baseline models across various logic reasoning tasks. Notably, compared with the best BoT, we observe significant accuracy improvements in 1-shot and 2-shot settings, particularly evidenced by increases of 14.7% and 15.3% on the Game of 24, 4.7% and 4.1% on Geometric Shapes, and 4.6% and 4.5% on Checkmate-in-One, demonstrating our flexibility and efficiency of leveraging cognitive preference in LLMs to activate logic capabilities through reverse reasoning, particularly with cognitive preference management.
Reasoning Efficiency.
Figure 4: Inference time comparison
Dataset RoT w/o PGRR w/o Logic w/o CPM
1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
Game of 24 97.5 97.8 77.5 77.8 86.0 86.5 92.0 93.5
Geometric Shapes 94.8 95.0 74.8 75.0 83.0 83.5 90.0 90.5
Multi-Step Arithmetic 98.6 99.0 78.6 79.0 87.5 88.0 95.0 95.5
WordSorting 99.8 100.0 79.1 81.0 89.0 89.2 98.0 98.5
Checkmate-in-One 91.8 92.0 80.8 82.0 81.0 83.5 98.0 88.5
$\times$×Python Puzzles 54.8 57.5 44.3 47.1 50.0 50.5 41.8 52.5
$\times$×MGSM (avg) 88.9 90.1 85.8 86.1 86.0 86.5 85.2 86.1
Sonnet Writing 89.5 92.0 72.5 73.0 87.0 77.5 89.0 91.5
As shown in Figure 4, RoT achieves competitive performance in reasoning efficiency, outperforming baselines and being second only to BoT, while BoT requires numerous initial golden thought templates in advance, which subtly limits its flexibility. In contrast, RoT focuses on exploring LLM-taste prompts, requiring only data for reasoning in subsequent steps after reverse reasoning warm-up.
5.3 Ablation Study
Impact of PGRR.
Excluding Preference-Guided Reverse Reasoning (PGRR) for exploring LLMs-taste prompts regarding tasks results in a significant reduction in overall task performance, with decreases of up to 20% in tasks like the Game of 24 (97.8% to 77.8%) and WordSorting (100% to 81%), indicating w/o PGRR weaken model’s cognition to task-specific requirements.
Impact of Logic.
Removing mathematical logic from RoT leads to notable declines in tasks requiring structured problem-solving, such as Multi-Step Arithmetic (99.0% to 88.0%) and Checkmate-in-One (92.0% to 83.5%), underscoring w/o Logic negatively affects structured problem-solving for complex reasoning.
Impact of CPM.
Lacking Cognitive Preference Manager (CPM) has a particularly pronounced effect on unknown tasks with Python Puzzles dropping from 57.5% to 52.5% and MGSM from 90.1% to 86.1%, indicating w/o CPM weaken RoT’s flexibility for tackling known tasks (case study for knwon task detailed in appendix B.1) and unknwon tasks (case study for unknwon task is detailed in Appendix B.2).
6 Conclusion
In this paper, we propose Reversal of Thought (RoT), a novel framework to enhance the logical reasoning capabilities of LLMs. By integrating reverse reasoning with meta-cognitive mechanisms and cognitive preference management, RoT improves reasoning accuracy and efficiency while minimizing computational costs, which leverages Preference-Guided Reverse Reasoning and Cognitive Preference Manager, which optimally aligns LLM reasoning processes with their cognitive preferences shaped by their pretraining and RLHF. Comprehensive experiments across diverse reasoning tasks demonstrate that RoT consistently outperforms state-of-the-art baselines in both known and unknown task scenarios, demonstrating the potential to expand knowledge boundaries through cognitive preference template. Our research provides valuable insights into future studies focused on further enhancing LLMs’ reasoning capacities by dynamically exploring cognitive preferences for complex reasoning tasks.
Limitations
In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
Ethical Considerations
Since the datasets used in our experiments focus on pure mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity, which are all sourced from publicly available datasets Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024) and devoid of any personal privacy or sensitive ethical information. Therefore, we do not identify any immediate ethical concerns regarding our current work. Additionally, we conduct human evaluations of cognitive preference consistency with three anonymous professional annotators following our instructions (detailed in appendix C), and they receive equivalent compensation.
Appendix A Prompt for CPM (Known/Unknown)
@@ B.1 Case Study for Known Task (Game of 24)
$\forall P(A,B,C,D)$ in permutations of [A, B, C, D]. $\forall op_{n}$ in (+, -, $\times$, $\div$) where $n\in\{0,1,2,3\}$. For each permute $\neg\exists(Expression==24)\lor(Expression=permute[0]\;op_{1}\;permute[1]\\ ;op_{2}\;permute[2]\;op_{3}\;permute[3])$: $\exists(Expression==24)\supset Print(Expression)\land Terminate\;Process.$ Print("No feasible solution exists."). Case Examples: Example 1: Input: 4 6 7 1 Output: One feasible solution (e.g., $6/(1-(4/7))=24$). Example 2: Input: 4 2 2 1 Output: "No feasible solution exists." Input-Output Format:
@@ B.2 Case Study for Unknown Task (MGSM)
(1) Style Consistency: How well responses maintain a consistent stylistic tone across different tasks. (2) Coherence: Internal logic and smooth flow of the response.

### v2

@@ 3. Formulate Logical Pseudocode:
A.3 Prompt for CPM (Known/Unknown)
@@ B.1.1 $P_{final}$ Case study for Game of 24
[…]
@@
B.1.2 Instantiation Case Study for Game of 24 Using $P_{final}$
Instantiation Case Study for Game of 24 Using $P_{final}$ Input: 4 9 10 13
Steps: Let me compute possible expressions for this input that result in 24 using provided logic pseudocode.
1. Generate all permutations of the numbers (4,9,10,13): This gives 24 possible arrangements.
2. Apply operations: For each permutation, test all combinations of the operations (+, -, *, /) between the numbers.
3. Check if the result equals 24: Evaluate the expression and check if it equals 24.
4. Output the result: If a valid expression is found, print it. If not, output that no feasible solution exists.
import itertools
inputs = [4,9,10,13]
operations = [’+’, ’-’, ’*’, ’/’]
def evaluate_expression(expr):
    try:
        return eval(expr)
    except ZeroDivisionError:
        return None
permutations = itertools.permutations(inputs)
op_combinations = itertools.product(operations, repeat=3)
solutions = []
for perm in permutations:
    for ops in op_combinations:
        exprs = [
            f"({perm[0]} {ops[0]} {perm[1]}) {ops[1]} ({perm[2]} {ops[2]} {perm[3]})",
            f"({perm[0]} {ops[0]} ({perm[1]} {ops[1]} {perm[2]})) {ops[2]} {perm[3]}",
            f"{perm[0]} {ops[0]} ({perm[1]} {ops[1]} ({perm[2]} {ops[2]} {perm[3]}))",
            f"({perm[0]} {ops[0]} {perm[1]}) {ops[1]} ({perm[2]} {ops[2]} {perm[3]})",
        for expr in exprs:
            if evaluate_expression(expr) == 24:
                solutions.append(expr)
print(solutions)
  Output: 4 * (9 + (10 - 13))=24
@@ B.2.1 $P_{final}$ Case study for MGSM
B.2.2 Instantiation Case Study for MGSM Using $P_{final}$
Instantiation Case Study for MGSM Using $P_{final}$ Input: Paul is at a train station and is waiting for his train. He isn’t sure how long he needs to wait, but he knows that the fourth train scheduled to arrive at the station is the one he needs to get on. The first train is scheduled to arrive in 10 minutes, and this train will stay in the station for 20 minutes. The second train is to arrive half an hour after the first train leaves the station, and this second train will stay in the station for a quarter of the amount of time that the first train stayed in the station. The third train is to arrive an hour after the second train leaves the station, and this third train is to leave the station immediately after it arrives. The fourth train will arrive 20 minutes after the third train leaves, and this is the train Paul will board. In total, how long, in minutes, will Paul wait for his train?
Steps:
Step 1: Symbol Information
1. $\alpha$ (first train arrival) = 10 minutes
2. $\beta$ (first train stay) = 20 minutes
3. $\gamma$ (second train stay as a quarter of the first train’s stay) = $\frac{\beta}{4}$
4. $\delta$ (time between second train departure and third train arrival) = 60 minutes
5. $\varepsilon$ (fourth train arrival after third train leaves) = 20 minutes
Step 2: Identify Contextual Clues and Operations
1. First train departure time = $\alpha+\beta$
2. Second train arrival time = First train departure time + 30 minutes
3. Second train stay time = $\gamma$
4. Third train arrival time = Second train departure time + $\delta$
5. Fourth train arrival time = Third train departure time + $\varepsilon$
Step 3: Perform Calculations Based on Identified Operations
[…]
2. Second train arrival time = Add(First train departure time,30) = $30+30=60$ minutes
[…]
4. Second train departure time = Add(Second train arrival time ,Second train stay time) = $60+5=65$ minutes
5. Third train arrival time = Add(Second train departure time,$\delta$) = $65+60=125$ minutes
6. Third train departure time = Third train arrival time (since it leaves immediately) = 125 minutes
7. Fourth train arrival time = Add(Third train departure time, $\varepsilon$) = $125+20=145$ minutes
Step 4: Calculate the Final Result
Final Result: Fourth train arrival time = 145 minutes
  Output: 145 minutes
@@ 7. Fourth train arrival time = Add(Third train departure time, $\varepsilon$) = $125+20=145$ minutes
(1) Style Consistency: How well responses maintain a consistent stylistic tone for thinking across different tasks. (2) Coherence: Internal logic and smooth flow of the response.